Thirty six bearings made by a certain process have a mean diameter of 0.64 cm and standard deviation 0.05 cm.

Construct:

  1. 99% confidence interval for the actual average diameter of bearings made by this process
  2. 88% confidence interval for the actual average diameter of bearings made by this process.
Answer #1

Solution :

Given that,

i

t\alpha /2,df = 2.724

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.724 * (0.05 / \sqrt 36)

Margin of error = E = 0.02

The 99% confidence interval estimate of the population mean is,

\bar x - E 9 \mu < \bar x + E

0.64 - 0.02 < \mu < 0.64 + 0.02

0.62 < \mu < 0.66

(0.62 , 0.66)

ii.

t\alpha /2,df = 1.594

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.594 * (0.05 / \sqrt 36)

Margin of error = E = 0.01

The 88% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

0.64 - 0.01 < \mu < 0.64 + 0.01

0.63 < \mu < 0.65

(0.63 , 0.65)